# Q-Mamba: On First Exploration of Vision Mamba for Image Quality Assessment

### v1

@@ 3.3 Tuning the Mamba with StylePrompt
Although the Mamba architecture offers reduced computational complexity compared to other models, achieving higher performance still necessitates a substantial number of parameters, which poses a challenge for efficient and transferable learning critical in Image Quality Assessment (IQA) tasks. Prompt tuning[19, 23], a common method for efficient fine-tuning, presents a promising approach. However, conventional prompt tuning techniques involving convolutional or attention operations still require significant parameter overhead[49]. Considering the need for transfer learning in IQA tasks to adapt specifically to the current distortion style, we propose the StylePrompt to enhance the transferability of IQA capabilities. Our method achieves results comparable to full-parameter fine-tuning with a minimal parameter increase.We illustrate the StylePrompt we used in Figure1. Our StylePrompt consists of two parts, which are described in detail below:
@@ 3.3.1 StylePrompt Generation (SPG)
In a multi-stage network architecture, as images progress through each stage of the network, we learn a set of prompts $P_{s}\in\mathbb{R}^{N\times 1\times 1\times{C}}$, containing N prompt components designed to generate affine parameters for style adaptation. These prompts are specifically utilized to inject the distortion style of the current image into the features $F_{i}\in\mathbb{R}^{\hat{H}\times\hat{W}\times\hat{C}}$, facilitating learning of the style pertinent to the current domain. We have devised a straightforward StylePrompt Generation stage where the prompt components $P_{s}$ are initially formed as a set of learnable parameters, interacting with the incoming features to embed distortion information.
To enable the prompt components to extract specific style information from the input stream dynamically, we predict the weights for different prompt components based on the input features. This process involves performing a global pooling on the current layer’s features followed by applying a softmax function to obtain the weights for the prompt group. These weights are then applied to the multiple prompt components to amalgamate them into a new prompt $P_{f}$, effectively encapsulating the current style information. The operation can be succinctly summarized by the following formula:
@@ 3.3.2 StylePrompt Injection (SPI)
In the StylePrompt Generation phase, we created the fused prompt $P_{f}$, which contains the distortion style information of the current domain. During the StylePrompt Injection process, this style information is injected into the original features $F_{i}\in\mathbb{R}^{\hat{H}\times\hat{W}\times\hat{C}}$ produced by the current layer. We adapt the distribution of the original features to the style by learning affine parameters. Specifically, as each layer utilizes a fixed-size prompt component, the fused Prompt generated in the StylePrompt Generation phase is matched to the channel dimensions of the current features through linear layers designed to align dimensions. This vector is then processed through two specific linear layers whose primary function is to adjust dimensions, thereby producing the affine parameters $\gamma_{v}\in\mathbb{R}^{1\times 1\times\hat{C}}$ and $\beta_{v}\in\mathbb{R}^{1\times 1\times\hat{C}}$. These parameters are used to inject the distorted style information solely on the channel dimension into the original feature distribution. The process of our StylePrompt Injection can be summarized as follows:
@@
$F{{}^{\prime}}_{i}=F_{i}\cdot(1+\gamma)+\beta$ (7)
@@
4 Experiments
4.1 Experimental Setup
4.1.1 Datasets
We conducted foundational experiments on ten popular IQA datasets, which include four synthetic datasets: LIVE[53], CSIQ[22], TID2013[48], and KADID[33]; four authentic datasets: LIVEC[10], KonIQ[17], LIVEFB[68], and SPAQ[6]; and two AIGC datasets: AIGC2023[61] and AGIQA3K[25]. In our single-task experiments, we utilized all ten datasets for testing. We report the results for the eight datasets consisting of synthetic and authentic datasets in the main text, including comparisons with existing SOTA BIQA methods and performance comparisons of Swin Transformer, ViT, and ResNet across different parameter versions. In the appendix, we report the comparative results of the two AIGC datasets against the aforementioned model architectures. In our universal IQA and transferable IQA experiments, we selected six datasets for training and testing, which include two synthetic datasets (LIVE[53], KADID[33]), two authentic datasets (LIVEC[10], KonIQ[17]), and two AIGC datasets (AIGC2023[61], AGIQA3K[25]).
4.1.2 Evaluation Criteria
4.1.3 Experimental Details
Our experimental methodology closely adheres to the training strategy outlined in DEIQT[50], where input images are randomly cropped into ten patches of 224×224 resolution. We utilized three variants of the VMamba architecture: VMamba-B, VMamba-S, and VMamba-T. Both VMamba-B and VMamba-S feature an encoder with a depth of 15 layers, where VMamba-B incorporates an embedding dimension of 128, and VMamba-S has an embedding dimension of 96. In contrast, VMamba-T is characterized by a reduced depth of 4 layers and an embedding dimension of 96. Training procedures leveraged weights pre-trained on the ImageNet-1K dataset, spanning across a total of 9 epochs, with batch sizes adjusted according to the respective dataset sizes. All experiments were conducted utilizing multiple NVIDIA RTX 4090 GPUs.
  Model LIVE CSIQ TID2013 KADID LIVEC KonIQ LIVEFB SPAQ
  FLOPs PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC Average
     ILNIQE[73] - 0.906 0.902 0.865 0.822 0.648 0.521 0.558 0.534 0.508 0.508 0.537 0.523 0.332 0.294 0.712 0.713 0.618
BRISQUE[45] - 0.944 0.929 0.748 0.812 0.571 0.626 0.567 0.528 0.629 0.629 0.685 0.681 0.341 0.303 0.817 0.809 0.664
WaDIQaM[2] - 0.955 0.960 0.844 0.852 0.855 0.835 0.752 0.739 0.671 0.682 0.807 0.804 0.467 0.455 - - 0.763
DBCNN[75] - 0.971 0.968 0.959 0.946 0.865 0.816 0.856 0.851 0.869 0.851 0.884 0.875 0.551 0.545 0.915 0.911 0.852
TIQA[70] - 0.965 0.949 0.838 0.825 0.858 0.846 0.855 0.850 0.861 0.845 0.903 0.892 0.581 0.541 - - 0.829
MetaIQA[79] - 0.959 0.960 0.908 0.899 0.868 0.856 0.775 0.762 0.802 0.835 0.856 0.887 0.507 0.540 - - 0.815
P2P-BM[69] - 0.958 0.959 0.902 0.899 0.856 0.862 0.849 0.840 0.842 0.844 0.885 0.872 0.598 0.526 - - 0.835
HyperIQA[56] - 0.966 0.962 0.942 0.923 0.858 0.840 0.845 0.852 0.882 0.859 0.917 0.906 0.602 0.544 0.915 0.911 0.858
TReS[11] - 0.968 0.969 0.942 0.922 0.883 0.863 0.858 0.859 0.877 0.846 0.928 0.915 0.625 0.554 - - 0.858
MUSIQ[21] - 0.911 0.940 0.893 0.871 0.815 0.773 0.872 0.875 0.746 0.702 0.928 0.916 0.661 0.566 0.921 0.918 0.832
DEIQT[50] - 0.982 0.980 0.963 0.946 0.908 0.892 0.887 0.889 0.894 0.875 0.934 0.921 0.663 0.571 0.923 0.919 0.884
LoDa[66] - 0.979 0.975 - - 0.901 0.869 0.936 0.931 0.899 0.876 0.944 0.932 0.679 0.578 0.928 0.925 0.882
@@ 4.2.1 Task-specific IQA
We conducted comprehensive training and testing across the ten datasets previously introduced, drawing analytical comparisons based on results reported by existing methods. Given the predominance of existing BIQA methods targeting synthetic and distorted datasets, we present in Table 1 a detailed comparative analysis of the performance of state-of-the-art (SOTA) methods alongside popular architectures such as ResNet[16], ViT[5], and Swin Transformer[40], against our Q-Mamba architecture. This comparison elucidates performance discrepancies and computational complexities across different parameter configurations. Our findings demonstrate that the LocalQMamba configuration achieves optimal performance across six datasets and, when juxtaposed with other models of comparable parameter scales, it exhibits lower Gflops, thereby confirming the efficacy and reduced computational complexity of the Mamba architecture. Comparative results for the two AIGC datasets are documented in the appendix in Table 7.
@@ 4.2.2 Universal IQA
The efficacy of the Q-Mamba architecture was rigorously assessed across a spectrum of datasets, adopting a joint training protocol on six varied datasets, encompassing two synthetic datasets (LIVE, KADID), two authentic datasets (LIVEC, KonIQ), and two AIGC datasets (AIGC2023, AGIQA3K). Subsequent to this joint training phase, we evaluated the performance outcomes, with the findings depicted in Table 2. These results highlight Q-Mamba’s robust multi-tasking capabilities. In comparison to several established foundational models, Q-Mamba consistently delivered notable performance across the majority of the datasets, thereby substantiating its versatility and effectiveness in managing generalized tasks with high efficacy.
Figure 3: t-SNE clustering of different backbones for various types of distortions
@@ 4.2.3 Analysis
To elucidate the reasons behind the VMamba architecture’s notable performance on task-specific IQA and universal IQA tasks, we engaged a subset of the KADID dataset featuring seven distinct types of distortions. We conducted t-SNE[59] dimensionality reduction clustering on the features processed through four disparate model architectures, as illustrated in Figure 3. The clustering outcomes demonstrate that the VMamba architecture exhibits exceptional distortion perception capabilities, effectively distinguishing various types of distortions and surpassing the performance of the ViT structure. Conversely, the ResNet and Swin Transformer architectures did not manifest significant abilities in aggregating different distortion types. Consequently, it becomes apparent that VMamba’s exemplary performance across diverse quality assessment tasks can be attributed, in part, to its robust distortion perception capabilities.
Table 2: Performance comparison for universal IQA. Bold indicates the top two results.

### v2

@@ 3.2 Tuning the Mamba with StylePrompt
Although the Mamba architecture reduces computational complexity compared to other models, achieving higher performance still requires a substantial number of parameters, which poses a challenge for efficient transferable learning, crucial for IQA tasks. We observe that domain shifts in IQA tasks tend to correlate with their feature statistics or style (Lu et al., 2022), such as mean and variance. Building on this insight, we propose a lightweight tuning strategy, StylePrompt, designed to adjust the mean and variance of the pre-trained QMamba features, thereby aligning them with the distortions and content types of the target domain. This approach enables us to achieve results comparable to full-parameter fine-tuning while using a minimal number of parameters, significantly enhancing both the efficiency and performance of QMamba in transferable IQA tasks. Figure 2 illustrates the StylePrompt, which consists of two components that will be described in detail below:
@@ 3.2.1 StylePrompt Generation (SPG)
We designed the StylePrompt Generation phase to facilitate the creation of prompts and their interaction with the original features. In a multi-stage network architecture, as images progress through each stage of the network, we learn a set of prompts $P_{s}\in\mathbb{R}^{N\!\times\!1\!\times\!1\!\times\!C}$, containing N prompt components designed to generate affine parameters for style adaptation. These prompts are specifically utilized to inject the distortion information of the current data into the features $F_{i}\in\mathbb{R}^{\hat{H}\times\hat{W}\times\hat{C}}$, facilitating learning of the style pertinent to the target domain.
To enable the prompt components to extract specific style information from the input stream dynamically, we predict the weights for different prompt components based on the input features. This process involves performing a global pooling on the current layer’s features followed by applying a softmax function to obtain the weights for the prompt group. These weights are then applied to the multiple prompt components to amalgamate them into a new prompt $P_{f}$, effectively encapsulating the current style information. The operation can be briefly summarized by the following formula:
@@ 3.2.2 StylePrompt Injection (SPI)
In the SPG phase, the fused prompt $P_{f}$ is created, containing the distortion style information of the target domain. During the subsequent StylePrompt Injection process, this style information is injected into the original features $F_{i}\in\mathbb{R}^{\hat{H}\times\hat{W}\times\hat{C}}$ produced by the current layer. To achieve this, $P_{f}$ is matched to the channel dimensions of the current features using linear layers designed for dimensional alignment. The processed prompt then generates affine parameters $\gamma_{v}\in\mathbb{R}^{1\times 1\times\hat{C}}$ and $\beta_{v}\in\mathbb{R}^{1\times 1\times\hat{C}}$, which modulate the mean and variance of the original feature distribution solely along the channel dimension. This adaptive adjustment ensures that the feature distribution is tailored to the distortion style of the target domain with minimal computational overhead.
The process of our StylePrompt Injection can be summarized as follows:
@@
$F{{}^{\prime}}_{i}=F_{i}\cdot(1+\gamma_{v})+\beta_{v}$ (4)
@@
  LIVE CSIQ TID2013 KADID LIVEC KonIQ LIVEFB SPAQ
    Method GFLOPS PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC Average
    ILNIQE - 0.906 0.902 0.865 0.822 0.648 0.521 0.558 0.534 0.508 0.508 0.537 0.523 0.332 0.294 0.712 0.713 0.618
BRISQUE - 0.944 0.929 0.748 0.812 0.571 0.626 0.567 0.528 0.629 0.629 0.685 0.681 0.341 0.303 0.817 0.809 0.664
WaDIQaM(5.24M) 2.43G 0.955 0.960 0.844 0.852 0.855 0.835 0.752 0.739 0.671 0.682 0.807 0.804 0.467 0.455 - - 0.763
DBCNN(15.31M) 16.51G 0.971 0.968 0.959 0.946 0.865 0.816 0.856 0.851 0.869 0.851 0.884 0.875 0.551 0.545 0.915 0.911 0.852
TIQA(23.68M) - 0.965 0.949 0.838 0.825 0.858 0.846 0.855 0.850 0.861 0.845 0.903 0.892 0.581 0.541 - - 0.829
MetaIQA(13.24M) 1.82G 0.959 0.960 0.908 0.899 0.868 0.856 0.775 0.762 0.802 0.835 0.856 0.887 0.507 0.540 - - 0.815
HyperIQA(27.38M) 4.31G 0.966 0.962 0.942 0.923 0.858 0.840 0.845 0.852 0.882 0.859 0.917 0.906 0.602 0.544 0.915 0.911 0.858
TReS(152.45M) 20.03G 0.968 0.969 0.942 0.922 0.883 0.863 0.858 0.859 0.877 0.846 0.928 0.915 0.625 0.554 - - 0.858
MUSIQ(27.13M) 9.02G 0.911 0.940 0.893 0.871 0.815 0.773 0.872 0.875 0.746 0.702 0.928 0.916 0.661 0.566 0.921 0.918 0.832
DEIQT(24.04M) 5.41G 0.982 0.980 0.963 0.946 0.908 0.892 0.887 0.889 0.894 0.875 0.934 0.921 0.663 0.571 0.923 0.919 0.884
LoDa(*) 23.74G 0.979 0.975 - - 0.901 0.869 0.936 0.931 0.899 0.876 0.944 0.932 0.679 0.578 0.928 0.925 0.882
@@
    LoDa has a total of 118.23M model parameters and 8.93M trainable parameters.
4 Experiments
4.1 Experimental Setup
4.1.1 Datasets
We conducted foundational experiments on ten popular IQA datasets, which include four synthetic datasets: LIVE(Sheikh et al., 2006), CSIQ(Larson & Chandler, 2010), TID2013(Ponomarenko et al., 2015), and KADID(Lin et al., 2019); four authentic datasets: LIVEC(Ghadiyaram & Bovik, 2015), KonIQ(Hosu et al., 2020), LIVEFB(Ying et al., 2020), and SPAQ(Fang et al., 2020); and two AIGC datasets: AIGC2023(Wang et al., 2023a) and AGIQA3K(Li et al., 2023a). We will provide a more detailed introduction to these datasets in Appendix.
4.1.2 Evaluation Criteria
4.1.3 Experimental Details
Our experimental methodology closely follows the training strategy outlined in DEIQT(Qin et al., 2023), where input images are randomly cropped into ten patches, each with a resolution of 224×224. We employed three variants of the VMamba architecture: QMamba-B, QMamba-S, and QMamba-T. Both QMamba-B and QMamba-S feature an encoder with a depth of 15 blocks, with QMamba-B incorporating an embedding dimension of 128, and QMamba-S using an embedding dimension of 96. In contrast, QMamba-T is designed with a reduced depth of 4 blocks and an embedding dimension of 96. Training procedures leveraged weights pre-trained on the ImageNet-1K dataset, spanning a total of 9 epochs. Batch sizes were adjusted according to the respective dataset sizes, e.g., 32 for LIVEC and 128 for KonIQ. We used the AdamW optimizer for training, with the learning rate set to $2\times 10^{-4}$ and a decay factor of 10 applied every 3 epochs. We compared the performance of ResNet(He et al., 2016), ViT(Dosovitskiy et al., 2020), and Swin Transformer(Liu et al., 2021), all of which were implemented using the official versions and loaded with pre-trained weights. To ensure fairness, other experimental settings were kept as consistent as possible. All experiments were conducted using multiple NVIDIA RTX 4090 GPUs.
@@ 4.2.1 Task-specific IQA
We conducted comprehensive training and testing across the ten datasets previously introduced, drawing analytical comparisons based on results reported by existing methods. For this task, $80\%$ of the images in each dataset were used for training, while the remaining $20\%$ were reserved for testing. Given the predominance of BIQA methods targeting synthetic and authentic datasets, Table 1 presents a detailed comparative analysis of state-of-the-art (SOTA) methods and popular architectures such as ResNet, ViT, and Swin Transformer, in comparison to our QMamba architecture. This comparison elucidates performance discrepancies and computational complexities across different parameter configurations. The results in the table demonstrate that the LQMamba configuration achieves optimal performance across six datasets. Compared to existing IQA methods with similar parameter counts, QMamba-T exhibits lower GFLOPS and superior performance, confirming the efficiency and reduced computational complexity of the Mamba architecture. Comparative results for the two AIGC datasets are documented in the appendix.
                 Mixed Training
                 LIVE & KADID & LIVEC & KonIQ & AGIQA3K & AIGCIQA2023
     Method Parameters GFLOPS PLCC_Average SRCC_Average
     ResNet-50 23.51M 4.11G 0.878 0.853
ViT-S 21.67M 4.61G 0.891 0.867
Swin-T 27.52M 4.51G 0.900 0.883
     DEIQT 24.04M 5.41G 0.895 0.873
LoDa 8.93M* 23.68G 0.876 0.855
     QMamba-T 27.99M 4.47G 0.905 0.886
LQMamba-T 29.87M 4.44G 0.909 0.888
  • *Trainable parameters.
Table 2: Performance Comparison for Universal IQA.
  Train KonIQ & LIVEC (Authentic) KADID & LIVE (Synthetic)
  Test KADID LIVE AIGC2023 AGIQA3K KonIQ LIVEC AIGC2023 AGIQA3K
    Fine-tuning Method PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC Average
    DEIQT 0.583 0.595 0.558 0.564 0.802 0.794 0.698 0.632 0.518 0.526 0.579 0.532 0.580 0.575 0.512 0.506 0.601
LoDa 0.600 0.600 0.733 0.729 0.808 0.800 0.715 0.650 0.527 0.513 0.558 0.499 0.592 0.586 0.564 0.541 0.622
Without_tuning 0.535 0.501 0.726 0.744 0.789 0.784 0.726 0.744 0.553 0.549 0.595 0.550 0.600 0.612 0.675 0.649 0.642
Lin_Probe(R) 0.606 0.581 0.783 0.811 0.824 0.804 0.769 0.695 0.871 0.848 0.785 0.751 0.803 0.788 0.771 0.723 0.755
Full_tuning (93.79M)
0.936 0.930 0.941 0.940 0.885 0.863 0.910 0.851 0.943 0.928 0.899 0.873 0.880 0.859 0.910 0.856 0.908
StylePrompt (3.83M)
0.920 0.912 0.949 0.948 0.877 0.854 0.908 0.854 0.932 0.913 0.888 0.866 0.880 0.865 0.906 0.852 0.901
StylePrompt & R 0.921 0.913 0.944 0.945 0.877 0.853 0.906 0.847 0.931 0.912 0.889 0.860 0.876 0.857 0.905 0.849 0.898
Table 3: Performance Comparison for Transferable IQA.
@@ 4.2.2 Universal IQA
We evaluated the effectiveness of the QMamba architecture for universal tasks by employing mixed training across six different datasets: two synthetic datasets (LIVE, KADID), two authentic datasets (LIVEC, KonIQ), and two AIGC datasets (AIGC2023, AGIQA3k). For each dataset, $20\%$ of the data was reserved for performance testing. The average results of models with similar scales are presented in Table 2, with detailed results provided in Appendix. These findings highlight strong multi-tasking capabilities of Mamba. Compared to several mainstream models, QMamba performed well across most datasets, confirming its effectiveness in handling general tasks.
@@ 4.2.3 Analysis
Our investigation into the efficacy of QMamba for IQA tasks reveals two key insights through analysis of the KADID dataset (7 distortion types) and cross-dataset validation. As shown in Figure 3, t-SNE visualization demonstrates superior distortion discrimination of QMamba: 1) Tightly clustered features for each distortion type, 2) Clear separation between dissimilar artifacts, and 3) Minimal inter-class overlap compared to ViT’s partial merging and CNN/Swin architectures’ significant feature entanglement.
This discriminative capability directly impacts practical performance. While QMamba demonstrates modest gains on simpler datasets such as LIVE and CSIQ, which contain only 4 to 5 distortion types, it achieves substantial improvements on more complex benchmarks like TID2013 and KADID, which include 24 to 25 distortion types. The architecture based on state-space modeling enables adaptive frequency processing through a selective scanning mechanism, dynamically emphasizing distortion-critical patterns while suppressing irrelevant features. This approach stands in contrast to convolutional networks with fixed receptive fields and Transformers that tend to over-mix local characteristics through global attention.
Our findings indicate that QMamba is well-suited for new quality assessment scenarios involving complex and mixed distortions, such as those introduced by neural compression or generative models. In these cases, traditional architectures often fail to capture subtle or entangled artifacts, while the selective modeling in QMamba provides better adaptability and robustness.
4.3 Efficient Transfer Learning for Mamba-Based IQA
In the context of transferable IQA tasks, we conducted domain-specific training using synthetic datasets (LIVE, KADID), authentic datasets (LIVEC, KonIQ), and AIGC datasets (AIGC2023, AGIQA3K). After training in one domain, models were directly transferred and tested on datasets from the other two domains. We employed the StylePrompt technique, as illustrated in Figure 2, where the architecture was kept intact by freezing all model parameters and fine-tuning only the StylePrompt module, which involved approximately $4\%$ of the total parameter count. This approach achieved performance levels comparable to those obtained through full-parameter training. The outcomes, presented in Table 3, clearly demonstrate the effectiveness and efficiency of the proposed StylePrompt method for transferable IQA tasks. The value of ”Average” represents the mean performance across all domain transfers, with more detailed results provided in the Appendix.
4.4 Ablation Study
4.4.1 Different Scanning Method
As shown in the task-specific results (Table 1) and universal evaluation results (Table 2), models adopting local scanning consistently outperform those based on cross-scanning strategies. This motivates our exploration of local scanning mechanisms, as employed in LQMamba.
Although the average performance gap between QMamba and LQMamba may appear small, further analysis reveals that LQMamba performs better on most individual datasets. The marginal overall gain is primarily due to relatively lower improvements on simpler datasets such as LIVE and CSIQ, which contain fewer distortion types and less diverse content. In contrast, on more challenging datasets like TID2013 and KADID-10k, which feature a wide variety of fine-grained distortions, LQMamba shows clear advantages (e.g., SRCC on TID2013: 0.964 vs. 0.949; on KADID: 0.941 vs. 0.932). These results highlight the effectiveness of the local scanning design in complex, distortion-rich scenarios where precise modeling of local artifacts is critical.
4.4.2 Different Model Scale
In our empirical analysis, as documented in Table 1, we discern that while QMamba-Base exhibits exceptionally robust quality perception capabilities, QMamba-Small either matches or exceeds the performance of QMamba-Base across the majority of the datasets. Although QMamba-Tiny displays a modest decline in performance metrics, it still delivers results that are competitive with current SOTA methods, solely utilizing the capabilities of QMamba-Tiny.
  Tuning Strategy Parameters PLCC_Avg. SRCC_Avg.
  SSF 6.1M 0.750 0.735
Crossattn_Prompt 12.17M 0.806 0.772
Conv_Prompt 28.33M 0.883 0.856
StylePrompt(ours) 3.83M 0.911 0.890
Table 4: Ablation Study on Different Prompt Tuning Strategies
4.4.3 Different Prompt Tuning Strategies
To validate the effectiveness of the StylePrompt Generation (SPG) process, we conducted tests by directly learning a set of affine parameters $\gamma$ and $\beta$ to modulate the original features, rather than using prompts for learning, similar to the SSF(Lian et al., 2022). Additionally, to assess the effectiveness of the StylePrompt interaction method, we explored various interaction strategies during tuning, including convolutional prompt interaction and cross-attention prompt interaction. The results in Table 4 demonstrate the superior efficiency and performance of StylePrompt.
In addition, we conducted several ablation studies related to the design of the prompts, including variations in the number and shape of the prompts. The detailed results of these experiments are provided in Appendix.
4.5 Discussion and Future Work
Looking ahead, the sequential modeling nature and efficiency of our SSM-based architecture make it a promising candidate for extension to video and audio quality assessment. Given the higher temporal complexity in VQA and the inherent sequential structure of audio signals, our method offers a unified and lightweight foundation for future research across visual and multimodal quality evaluation tasks.
5 Conclusion
In this paper, we introduced QMamba, a novel state space framework for image quality assessment that integrates task-specific evaluation, universal perception, and cross-domain transferability. Extensive experiments demonstrate that QMamba consistently outperforms established vision models, achieving significant improvements in accuracy while requiring substantially lower computational resources. The proposed StylePrompt mechanism enables robust cross-domain adaptation through lightweight dynamic feature recalibration, setting new benchmarks for IQA and demonstrating strong practical potential in real-world applications.
Appendix A Appendix / supplemental material
A.1 Detailed Description of the Datasets
We evaluate the performance of the proposed method across eight widely recognized BIQA datasets, comprising both synthetic and authentic datasets. The synthetic datasets include LIVE(Sheikh et al., 2006), CSIQ(Larson & Chandler, 2010), TID2013(Ponomarenko et al., 2015), and KADID-10k(Lin et al., 2019). These datasets feature a small number of pristine images that are synthetically distorted using various techniques such as JPEG compression and Gaussian blurring. Specifically, LIVE contains 799 images with 5 types of distortion, CSIQ includes 866 images with 6 distortion types, TID2013 comprises 3000 images with 24 distortion types, and KADID-10k includes 10125 images with 25 distortion types.
On the other hand, the authentic datasets include LIVEC(Ghadiyaram & Bovik, 2015), KonIQ-10k(Hosu et al., 2020), SPAQ(Fang et al., 2020), and FLIVE(Ying et al., 2020). LIVEC consists of 1,162 images with diverse authentic distortions captured by mobile devices. KonIQ-10k is composed of 10,073 images selected from YFCC-100M, covering a wide range of distortions such as brightness, colorfulness, contrast, noise, and sharpness. SPAQ contains 11,125 images collected by various smartphones, representing a large variety of scene categories. FLIVE, the largest in-the-wild IQA dataset to date, contains 39,810 real-world images with diverse content, sizes, and aspect ratios.
In response to the rapid development of AI-generated content, we also employed two additional datasets: AIGCIQA2023(Wang et al., 2023a) and AGIQA3K(Li et al., 2023a). AIGCIQA2023 contains over 2000 images generated by six state-of-the-art text-to-image models, evaluated through a subjective experiment on quality, authenticity, and correspondence. AGIQA3K consists of 2,982 images from GAN, autoregression, and diffusion-based models, with annotations for perceptual quality and text-to-image alignment.
A.2 Additional Experimental Results
A.2.1 Task-specific IQA results on AIGC datasets
  Model AGIQA3K AIGCIQA2023
  GFLOPS PLCC SRCC PLCC SRCC Average
     ResNet-50(23.51M) 4.11G 0.901 0.840 0.795 0.797 0.833
ResNet-101(42.50M) 7.83G 0.907 0.847 0.834 0.831 0.855
ResNet-152(58.15M) 11.53G 0.901 0.832 0.841 0.834 0.852
ViT-T(5.52M) 1.26G 0.865 0.787 0.760 0.766 0.795
ViT-S(21.67M) 4.61G 0.891 0.819 0.842 0.822 0.844
ViT-B(85.80M) 17.58G 0.897 0.830 0.853 0.835 0.854
Swin-T(27.52M) 4.51G 0.906 0.847 0.867 0.844 0.866
Swin-S(48.84M) 8.77G 0.908 0.849 0.875 0.857 0.872
Swin-B(86.74M) 15.47G 0.909 0.852 0.886 0.863 0.878
QMamba-T (27.99M) 4.47G 0.913 0.858 0.888 0.873 0.883
QMamba-S (49.37M) 8.71G 0.912 0.858 0.889 0.875 0.884
QMamba-B (87.53M) 15.35G 0.914 0.861 0.886 0.868 0.882
LQMamba-T(29.87M) 4.44G 0.914 0.862 0.884 0.868 0.882
LQMamba-S(52.91M) 8.66G 0.913 0.864 0.888 0.869 0.884
LQMamba-B(93.79M) 15.30G 0.915 0.858 0.888 0.871 0.883
Table 5: Results of AIGC for task-specific IQA on various backbones
A.2.2 Universal IQA detailed data
We provide the detailed results of Universal IQA in Table 6.